OpenStudy (anonymous):

A cube with edge length x and a right cylinder with circular base have the same volume. Find x ,in cm, if the radius of the cylinder is 5 cm and its height is 5 cm.

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

Setup of the problem is important. The volume of the cube is \[x^{3}\] We set that equal to the volume of the cylinder. The volume of a right cylinder is: \[base area * height\] So we have: \[x ^{3} = \pi(5^{2}cm)(5cm)\] \[x ^{3}= \pi125cm ^{3}\] Take the 1/3 power on both sides and you get: \[x = 5\sqrt[3]{\pi}cm\]

OpenStudy (anonymous):

it is the same 5 (pi)1/3?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok I understand now. Thanks for your help.

OpenStudy (anonymous):

The 3 in front is a fractional power. Just as square root would be x^(1/2)